OpenStudy (anonymous):

What are the steps you would use to convert 5.8 meters per hour to the unit cm/min?

OpenStudy (anonymous):

s this correct 5.8 m/hr cm/min 1m=100cm 1hr=60min 5.8*(1m/1hr)=5.8*(100cm/60min) 5.8 m/hr=9.667 cm/min

thomaster (thomaster):

yea just multiply by 100 to get 580 cm/hour then divide by 60 to get 9.667 cm/min \(\huge\hspace{200pt}\ddot\smile\)

OpenStudy (anonymous):

thank you